If 3.01×10 20 molecules are removed from 98mg of H 2 SO 4 , then number of moles of H 2 SO 4 left will be?

If 3.01×1020 molecules are removed from 98mg of H2SO4, then number of moles of H2SO4 left will be?

Grade:12

1 Answers

Arun
6 years ago

3.01 * 1020molecules = 3.01 * 1020/6.02 * 1023= 2 * 10-3moles

98 mg of H2SO4= 98 * 10-2/98 = 0.01 moles of H2SO4

Thus when 0.002 moles are removed from 0.01 moles no. of moles left = 0.008 moles